# MEMERAG: A Multilingual End-to-End Meta-Evaluation Benchmark for Retrieval Augmented Generation

### v3

@@ Abstract
María Andrea Cruz Blandón1thanks: Work done as an intern at Amazon, Jayasimha Talur2thanks: Correspondence: talurj@amazon.com, Bruno Charron2, Dong Liu2, Saab Mansour2, Marcello Federico2 1Tampere University, 2Amazon
@@ 1 Introduction
In a reference-free evaluation setup, gold multidimensional judgements (factuality, relevance, etc) of model generations can be leveraged in a meta-evaluation framework. In this framework, automated evaluators are evaluated against the human judgement to measure correlation. The automated evaluators can then be applied to measure the performance of new models’ outputs.
Previous work for RAG meta-evaluation mainly focused on English Fan et al. (2024) or leveraged human or machine translation of English datasets Sharma et al. (2024). Multilingual meta-evaluation is important to reliably measure performance across languages which can vary depending on language characteristics (low vs. high resource, complex morphology, etc) and scripts (Latin vs. non-Latin). Translation-based benchmarks, while permitting cross-language comparisons, suffer from translationese phenomena such as introducing simpler syntax and lexical choices Baker et al. (1993); Graham et al. (2020), thus leading to data distributionally different from native data and not necessarily reflecting native users preferences Chen et al. (2024). Our position is that translation-based (parallel) benchmarks should be complemented by native multilingual benchmarks.
@@ 3.1 Question Selection
The questions in the MIRACL dataset were generated by humans based on prompts. This leads to questions that may be answerable by the prompt but may have ambiguity outside of that context. In particular, we identified as problematic the questions for which the right answer can change over time. For example “Who is the president of Spain?” or “How old is Drake Hogestyn?”. In a RAG setting, different passages may have been written at different times and provide conflicting context. Additionally, the time of reference is usually not explicit in the question. To remove those complications we automatically filtered out time-dependent questions across all languages444 See Appendix G for all prompts used during dataset construction. . We combined the train and dev splits of the MIRACL dataset, which corresponds to a total of 3,662, 305, 2,810, 1,486 and 1,519 questions respectively for EN, DE, ES, FR and HI. Of those, 244, 13, 120, 64 and 86 were identified as time-dependent and filtered out.
@@ 3.3 Answer Generation
After question and passage selection, we generate an answer for each question-context pair and each of five state-of-the-art LLMs. [^5] We generated answers with Claude 3 Sonnet, Llama3 70B, Llama3 8B, Mistral 7B, and GPT-4o mini. Those LLMs were selected to cover a range of model sizes, open weight and proprietary models. We prompted all the models in English [^6] There is evidence in the literature for better model accuracy when the models are prompted to ”think” in English Lai et al. (2023); Liu et al. (2024), though under particular scenarios other languages could perform better Behzad et al. (2024). We leave additional prompting experiments for future work., asking to answer the question based only on the given context, and requesting the answer to be provided in the same language as the context and question. For all models, we set the temperature to $0.1$, and maximum number of output tokens to 1000.
@@ 3.5 Annotation Process
From the question-context-answer triplets obtained in Section 3.3, we randomly sampled 250 questions per language (50 per answer-generating model, without overlapping questions for diversity). We employed a professional vendor with native annotators [^7] Annotators were compensated with a competitive hourly rate that is benchmarked against similar roles in their country of residence. to gather annotations for each sentence  [^8] Sentences were segmented using the pySBDSadvilkar and Neumann (2020) package. of the 250 answers per language. Among the 250 answers per language, a random subset of 10 were assigned to 3 annotators for computing the IAA and the rest to a single annotator. The statistics of the annotated dataset are presented in Table 8.
The annotations were gathered via a web-based tool that implemented the flow chart of the annotation guidelines (see Appendix A for details). To further enhance IAA, we drew upon the findings of Krishna et al. (2023), which demonstrated that highlighting relevant information aids annotators in performing tasks and reaching consensus. Thus, we utilized the Llama 3 70B LLM to identify sentences within the retrieved passages that could potentially serve as supporting information to the answer sentences.
@@
Table 2 summarizes the IAA per language for faithfulness and relevance labels assigned by 3 annotators. We report IAA using Gwet’s AC1 Gwet (2008) and Fleiss Kappa Fleiss (1971). We observe high agreement for faithfulness (0.84-0.93 Gwet’s AC1 and 0.70-0.88 Fleiss Kappa) and even higher agreement for relevance (0.95-1.0 Gwet’s AC1 and 0.63-1.0 Fleiss Kappa). This shows that the annotators are aligned and indicates a high quality of the annotations. Comparing to previous work, Tang et al. (2024b) report a Fleiss Kappa of 0.34-0.42 on faithfulness labels which they deem fair to moderate agreement. Note that a direct comparison with this work is not possible as they deal with different tasks, nevertheless the high IAA we are reporting is a testament of the effectiveness of our flow chart-based annotation design. Further details on IAA including fine-grained explanatory labels can be found in Appendix B.
@@ 4 Annotation Results
Table 3 shows the distribution of faithfulness and relevance labels across the five languages. The distribution of labels is consistent across all languages, with a few exceptions. On faithfulness, German and Hindi show higher percentages of Supported answers. On relevance, Spanish presents a significantly higher percentage of labels Adds context to the answer compared to other languages while English had a very low share of labels Unrelated to the question. As a partial explanation for this, we note that Spanish questions generated the largest numbers of output sentences, 563 vs. 400 (see Table 8) for English. Hence, we expect that the relevance statistics reflect the tendency of Spanish answers to be more verbose.
@@
For a more granular insights into the annotations, we show in Table 4 the distribution of the explanatory fine-grained faithfulness labels for each language. (As label names are quite self-explanatory, we refer for their precise meaning to Figure 3 in Appendix A.) Table 4 is split into three blocks, respectively, addressing fine-grained labels for Supported (top), Not supported (middle), and Challenging to determine (bottom) answers. We observe significant differences across languages. For example, supported answers for English are prominently under form of a logical conclusion from the context (41.2%), for German and Hindi from direct paraphrasing of information in the context (41.7% and 40.7%), while for French and Spanish from a more balanced combination of the two reasons. On the side of unsupported answers, the main mistake type in English is Wrong reasoning (10%), i.e. answers are non logical conclusions from the context, while this type or error is significantly rarer (under 2%) for all other languages. The rate of Adds new information errors, a.k.a. hallucinations, ranges from 7% for English to 16% for Spanish. The observed cross-linguistic variations in the distribution of labels can be attributed to the different nature of the questions and accuracy of the models across the 5 languages.
@@ 5 Dataset Applications
The MEMERAG dataset is designed to support the development of reliable automatic evaluation methods. For that purpose, we describe in this section how our dataset can be used as a benchmark to enable various meta-evaluation use cases. We focus on two applications: 1) Prompt selection: The ability of our benchmark to effectively select prompts for automatic evaluation, 2) Model selection: The effectiveness of our benchmark to distinguish and select models. By concentrating on these aspects, we can evaluate the benchmark’s utility as a comprehensive tool for multilingual model assessment. While we provide reference baselines for each application, our focus is on showcasing the effectiveness of the benchmark rather than the underlying capabilities of the LLMs.
@@ 5.1 Experimental Setup
Our benchmark is composed of multiple tasks defined by the annotation dimension and subset considered. On the annotation dimension, we focus our experiments on the coarse-grained faithfulness dimension. This dimension is more challenging than relevance as highlighted by the lower IAA, while retaining a high-enough IAA to make for a trustworthy benchmark. We invite benchmark users to also experiment on the other dimensions provided by the dataset depending on their use case. Note that we remove the sentences labelled as Challenging to determine by human annotators. On the annotation subset, we first distinguish the multilingual task which uses the full dataset and the monolingual task, which only considers a single language. Those task can then be further broken down at the fine-grained level by considering the subset of sentences with a certain fine-grained label. Performance is evaluated with Balanced Accuracy (BAcc) (see Appendix F for definition), with equal weights on each coarse-grained label and language. We conduct significance testing using permutation tests Good (2013), with further details in Appendix E.
@@
To demonstrate how the benchmark can be used to select the appropriate prompt, we experiment with multiple prompting strategies from simple to advanced techniques, starting with zero-shot prompting, where LLMs directly classify statements as Supported or Not supported. We then implement chain-of-thought (COT) prompting Wei et al. (2022), which incorporates an intermediate reasoning step. While these basic prompting strategies provide a good starting point for evaluation, our initial experiments revealed limitations in their ability to capture the nuanced requirements of our specific task. Without explicit guidelines in the prompt, automatic evaluators rely on their “world knowledge”, which may not align with the specific requirements of the evaluation task. To overcome this, we add instructions from the annotation guidelines (AG) in the prompt. Adding annotation guidelines provides clear criteria for what constitutes Supported versus Not Supported sentences, reducing ambiguity in the evaluation process. The various prompts are presented in Appendix H.
@@ 5.2 Experimental Results
Figure  2, shows the performance per language of various automatic evaluators with a fixed prompt (AG + COT), which allows us to select the best model for each language. We observe that GPT-4o mini performs best in English. For the rest of the languages Qwen 2.5 32B performs the best however, the results are not statistically different from GPT-4o mini. Our benchmark also provides users with the capability to conduct detailed, fine-grained analyses of model performance across various dimensions of faithfulness. The breakdown of automatic evaluation performance by error type is shown in Appendix J.
@@ Appendix E Statistical Significance Test Details
Appendix F Balanced Accuracy
Due to the presence of class imbalance in our dataset, we employ balanced accuracy as the primary metric to assess the performance of our automatic evaluators. Balanced accuracy provides a more robust measure of performance when dealing with imbalanced classes, as it equally weights the recall of each class.
For our binary classification task, where the automatic evaluators predict either "Supported" or "Not Supported", the balanced accuracy (BAcc) is calculated as follows:
$BAcc=\frac{1}{2}\left(\frac{TP}{TP+FN}+\frac{TN}{TN+FP}\right)$
Where TP, TN, FN, and FP denote True Positives (correctly identified “Supported” instances), True Negatives (correctly identified “Not Supported” instances), False Negatives (“Supported” instances misclassified as “Not Supported”), and False Positives (“Not Supported” instances misclassified as “Supported”), respectively.
Appendix G Prompts used for dataset construction
In section we describe the various prompts used in our pipeline. Out prompts are written as Jinja2101010https://jinja.palletsprojects.com/en/stable/ templates.
G.1 Time-dependent answer filtering
@@
G.2 Highlighting relevant segments
@@
G.3 Answer generation prompt
@@
Appendix H Prompts for automated evaluation
@@
Appendix I More Results on Meta-Evaluators
@@
Appendix J Fine grained Analysis

### v4

@@ Abstract
María Andrea Cruz Blandón1thanks: Work done by first author during an internship at Amazon. First two authors contributed equally to this work. Corresponding author: Jayasimha Talur email talurj@amazon.com., Jayasimha Talur Bruno Charron, Dong Liu, Saab Mansour, Marcello Federico 1Tampere University, Amazon
@@ 1 Introduction
In a reference-free evaluation setup, gold multidimensional judgements (factuality, relevance, etc) of model generations can be leveraged in a meta-evaluation framework. In this framework, automated evaluators are evaluated against human judgement to measure correlation. The automated evaluators can then be applied to measure the performance of new models’ output.
Previous work for RAG meta-evaluation mainly focused on English Fan et al. (2024) or leveraged human or machine translation of English datasets Sharma et al. (2024). Multilingual meta-evaluation is important to reliably measure performance across languages which can vary depending on language characteristics (low vs. high resource, complex morphology, etc.) and scripts (Latin vs. non-Latin). Translation-based benchmarks, while permitting cross-language comparisons, suffer from translationese phenomena such as introducing simpler syntax and lexical choices Baker et al. (1993); Graham et al. (2020), thus leading to data distributionally different from native data and not necessarily reflecting native users preferences Chen et al. (2024). Our position is that translation-based (parallel) benchmarks should be complemented by native multilingual benchmarks.
@@ 3.1 Question Selection
The questions in the MIRACL dataset were generated by humans based on prompts. This leads to questions that may be answerable by the prompt but may have ambiguity outside of that context. In particular, we identified as problematic the questions for which the right answer can change over time. For example “Who is the president of Spain?” or “How old is Drake Hogestyn?”. In a RAG setting, different passages may have been written at different times and provide conflicting context. Additionally, the time of reference is usually not explicit in the question. To remove those complications we automatically filtered out time-dependent questions across all languages444 See Appendix F for all prompts used during dataset construction. . We combined the train and dev splits of the MIRACL dataset, which corresponds to a total of 3,662, 305, 2,810, 1,486 and 1,519 questions respectively for EN, DE, ES, FR and HI. Of those, 244, 13, 120, 64 and 86 were identified as time-dependent and filtered out.
@@ 3.3 Answer Generation
After question and passage selection, we generate an answer for each question-context pair and each of five state-of-the-art LLMs. We generated answers with Claude 3 Sonnet, Llama3 70B, Llama3 8B, Mistral 7B, and GPT-4o mini. Those LLMs were selected to cover a range of model sizes, open weight and proprietary models. We prompted all the models in English [^5] There is evidence in the literature for better model accuracy when the models are prompted to process in English Lai et al. (2023); Liu et al. (2024), though under particular scenarios other languages could perform better Behzad et al. (2024). We leave additional prompting experiments for future work., asking to answer the question based only on the given context, and requesting the answer to be provided in the same language as the context and question. For all models, we set the temperature to $0.1$, and maximum number of output tokens to 1000.
@@ 3.5 Annotation Process
From the question-context-answer triplets obtained in Section 3.3, we randomly sampled 250 questions per language (50 per answer-generating model, without overlapping questions for diversity). We employed a professional vendor with native annotators [^6] Annotators were compensated with a competitive hourly rate that is benchmarked against similar roles in their country of residence. to gather annotations for each sentence  [^7] Sentences were segmented using the pySBDSadvilkar and Neumann (2020) package. of the 250 answers per language. The statistics of the annotated dataset are presented in Table 1. Among the 250 answers per language, a random subset of 10 were assigned to 3 annotators for computing the IAA and the rest to a single annotator (see Table 2).
The annotations were gathered via a web-based tool that implemented the flow chart of the annotation guidelines (see Appendix A for details). To further enhance IAA, we drew upon the findings of Krishna et al. (2023), which demonstrated that highlighting relevant information aids annotators in performing tasks and reaching consensus. Thus, we utilized the Llama 3 70B LLM to identify sentences within the retrieved passages that could potentially serve as supporting information to the answer sentences (see Appendix F).
@@
Table 2 summarizes the IAA per language for faithfulness and relevance labels assigned by 3 annotators. We report IAA using Gwet’s AC1 Gwet (2008) and Fleiss Kappa Fleiss (1971). We observe high agreement for faithfulness (0.84-0.93 Gwet’s AC1 and 0.70-0.88 Fleiss Kappa) and even higher agreement for relevance (0.95-1.0 Gwet’s AC1 and 0.63-1.0 Fleiss Kappa). This shows that the annotators are aligned and indicates a high quality of the annotations. Comparing to previous work, Tang et al. (2024b) report a Fleiss Kappa of 0.34-0.42 on faithfulness labels which they deem fair to moderate agreement. Note that a direct comparison with this work is not possible as they deal with different tasks, nevertheless the high IAA we are reporting is a testament of the effectiveness of our flow chart-based annotation design. Further details on IAA with fine-grained explanatory labels, and an extended version of the dataset (MEMERAG-Ext) with five annotators for 150 answers per language can be found in Appendix B.
@@ 4 Annotation Results
Table 3 shows the distribution of faithfulness and relevance labels across the five languages. The distribution of labels is consistent across all languages, with a few exceptions. On faithfulness, German and Hindi show higher percentages of Supported answers. On relevance, Spanish presents a significantly higher percentage of labels Adds context to the answer compared to other languages while English had a very low share of labels Unrelated to the question. As a partial explanation for this, we note that Spanish questions generated the largest numbers of output sentences with 52.1 words vs., for example, 30.3 words for English (see Table 1). Hence, we expect that the relevance statistics reflect the tendency of Spanish answers to be more verbose.
@@
For a more granular insights into the annotations, we show in Table 4 the distribution of the explanatory fine-grained faithfulness labels for each language. As label names are quite self-explanatory, we refer for their precise meaning to Figure 3 in Appendix A. Table 4 is split into three blocks, respectively, addressing fine-grained labels for Supported (top), Not supported (middle), and Challenging to determine (bottom) answers. We observe remarkable differences across languages. For example, supported answers for English are prominently under form of a logical conclusion from the context (41.2%), for German and Hindi from direct paraphrasing of information in the context (41.7% and 40.7%), while for French and Spanish from a more balanced combination of the two reasons. On the side of unsupported answers, the main mistake type in English is Wrong reasoning (10%), i.e. answers are non logical conclusions from the context, while this type or error is significantly rarer (under 2%) for all other languages. The rate of Adds new information errors, a.k.a. hallucinations, ranges from 7% for English to 16% for Spanish. The observed cross-linguistic variations in the distribution of labels can be attributed to the different nature of the questions and accuracy of the models across the 5 languages.
@@ 5 Dataset Applications
The MEMERAG dataset is designed to support the development of reliable automatic evaluation methods. For that purpose, we describe in this section how our dataset can be used as a benchmark to enable various meta-evaluation use cases. We focus on two applications: 1) Prompt selection: The ability of our benchmark to effectively select prompts for automatic evaluation, 2) Model selection: The effectiveness of our benchmark to compare and select models. By concentrating on these aspects, we can evaluate the benchmark’s utility as a comprehensive tool for multilingual model assessment. While we provide reference baselines for each application, our focus is on showcasing the effectiveness of the benchmark rather than the underlying capabilities of the LLMs.
@@ 5.1 Experimental Setup
Our benchmark is composed of multiple tasks defined by the annotation dimension and subset considered. On the annotation dimension, we focus our experiments on the coarse-grained faithfulness dimension. This dimension is more challenging than relevance as highlighted by the lower IAA, while retaining a high-enough IAA to make for a trustworthy benchmark. We invite benchmark users to also experiment on the other dimensions provided by the dataset depending on their use case. Note that for this experiment, we remove the sentences labelled as Challenging to determine by human annotators. On the annotation subset, we first distinguish the multilingual task which uses the full dataset and the monolingual task, which only considers a single language. Those tasks can then be further broken down at the fine-grained level by considering the subset of sentences with a certain fine-grained label. Performance is evaluated with Balanced Accuracy (BAcc), with equal weights on each coarse-grained label and language. We conduct significance testing using permutation tests Good (2013), with further details in Appendix E.
@@
To demonstrate how the benchmark can be used to select the appropriate prompt, we experiment with multiple prompting strategies from simple to advanced techniques, starting with zero-shot prompting, where LLMs directly classify statements as Supported or Not supported. We then implement chain-of-thought (COT) prompting Wei et al. (2022), which incorporates an intermediate reasoning step. While these basic prompting strategies provide a good starting point for evaluation, our initial experiments revealed limitations in their ability to capture the nuanced requirements of our specific task. Without explicit guidelines in the prompt, automatic evaluators rely on their “world knowledge”, which may not align with the specific requirements of the evaluation task. To overcome this, we add instructions from the annotation guidelines (AG) in the prompt. Adding annotation guidelines provides clear criteria for what constitutes Supported versus Not Supported sentences, reducing ambiguity in the evaluation process. The various prompts are presented in Appendix G.
@@ 5.2 Experimental Results
Figure  2, shows the performance per language of various automatic evaluators with a fixed prompt (AG + COT), which allows us to select the best model for each language. We observe that GPT-4o mini performs best in English. For the rest of the languages Qwen 2.5 32B performs the best however, the results are not statistically different from GPT-4o mini. Our benchmark also provides users with the capability to conduct detailed, fine-grained analyses of model performance across various dimensions of faithfulness. The breakdown of automatic evaluation performance by error type is shown in Appendix I.
@@ Appendix E Statistical Significance Test Details
Appendix F Prompts used for dataset construction
In section we describe the various prompts used in our pipeline. Out prompts are written as Jinja2999https://jinja.palletsprojects.com/en/stable/ templates.
F.1 Time-dependent answer filtering
@@
F.2 Highlighting relevant segments
@@
F.3 Answer generation prompt
@@
Appendix G Prompts for automated evaluation
@@
Appendix H More Results on Meta-Evaluators
@@
Appendix I Fine grained Analysis